### v1

# GeMID: Generalizable Models for IoT Device Identification

Abstract
With the proliferation of Internet of Things (IoT) devices, ensuring their security has become paramount. Device identification (DI), which distinguishes IoT devices based on their traffic patterns, plays a crucial role in both differentiating devices and identifying vulnerable ones, closing a serious security gap. However, existing approaches to DI that build machine learning models often overlook the challenge of model generalizability across diverse network environments. In this study, we propose a novel framework to address this limitation and evaluate the generalizability of DI models across datasets collected within different network environments. Our approach involves a two-step process: first, we develop a feature and model selection method that is more robust to generalization issues by using a genetic algorithm with external feedback and datasets from distinct environments to refine the selections. Second, the resulting DI models are then tested on further independent datasets in order to robustly assess their generalizability. We demonstrate the effectiveness of our method by empirically comparing it to alternatives, highlighting how fundamental limitations of commonly employed techniques such as sliding window and flow statistics limit their generalizability. Our findings advance research in IoT security and device identification, offering insights into improving model effectiveness and mitigating risks in IoT networks.
Index Terms:
IoT security, device identification, machine learning, generalizability.
I Introduction
The Internet of Things (IoT) seamlessly integrates our cyber world with the physical world and is becoming increasingly prevalent in daily life. According to published statistics, the number of IoT devices has exceeded 15 billion and is projected to reach approximately 30 billion by 2030 [1].
Despite the rapid proliferation of devices and providers, security remains a critical issue. IoT devices can be more challenging to secure than conventional devices due to their diverse hardware, software, and varied manufacturer profiles [2]. Moreover, unlike conventional devices, many IoT devices have non-standard interfaces which impair user interaction and make it difficult to mitigate against vulnerabilities [3]. The NETSCOUT Threat Intelligence Report [4] indicates that a new IoT device on the network typically faces its first attack within 5 hours and becomes a specific attack target within 24 hours, and that most attacks exploit vulnerabilities in IoT devices [5]. Beyond being targets in attacks, compromised IoT devices can serve as tools in botnet attacks, exemplified by Mirai, URSNIF, and BASHLITE [6], where captured devices are used to orchestrate high-volume Distributed Denial of Service (DDoS) attacks. While various methods have been developed to detect and mitigate ongoing attacks, primarily Intrusion Detection Systems (IDS), proactive measures such as device-specific updates, internet access restrictions or isolation can prevent these attacks before they happen. Given the variety, quantity and unfamiliar interfaces of IoT devices, it is impractical to rely on users to implement these solutions. As a result, DI systems have been developed to automatically identify devices based on their activity and make it possible to apply appropriate security measures to create a secure IoT ecosystem.
Most existing approaches to DI involve the construction of Machine Learning (ML) models. When developing ML models, it is important to ensure that they generalize beyond the specific environment in which they were trained. In DI, this means that models trained on data from one device should be able to detect other devices of the same make and model operating within different network environments. Achieving this requires accounting for the domain shifts and variations typical of network and IoT environments [7]. In this regard, a prominent limitation of previous work is that most studies have relied upon a single dataset collected in a single network environment for both developing and testing their models [8, 9, 10]. This has resulted in an incomplete, and potentially misleading, picture of DI generalizability. Where multiple datasets have been used, the focus has been on the generalizability of methodologies and feature sets rather than of model instances [11, 12, 13].
In this study, we explicitly focus on developing DI model instances that are demonstrably generalizable across network environments. We measure this using multiple datasets containing the same devices operating in different network environments. We use a two-stage process to develop the models. The first stage involves feature and model selection, and focuses on identifying device-independent features that do not display over-dependence on their network environment. In the second stage, we use an independent dataset to train device-specific model instances. We then use a further dataset, containing the same devices operating in a different network environment, to evaluate the generalizability of these device-specific models.
We make the following contributions to the field of DI:
  1. 1.
    Comprehensive evaluation of model generalizability: We provide an extensive assessment of ML-based DI models across diverse network environments, highlighting key factors that influence generalizability.
  2. 2.
    Novel study framework for assessing generalizability: We introduce a two stage framework specifically designed to evaluate the generalizability of DI models across varied conditions, offering a more rigorous methodology for testing model robustness.
  3. 3.
    Insight into feature selection and its impact on generalizability: We demonstrate how the method of feature selection and construction critically impacts the generalizability of DI models. In particular, we show that features derived from individual packet characteristics lead to superior generalizability compared to those based on flow or window statistics.
  4. 4.
    Validation of packet-based approaches: Through empirical comparisons, we validate our packet-based method against other approaches, demonstrating that models built on individual packet features consistently outperform methods that use flow or window-based statistics in terms of generalizability.
  5. 5.
    Commitment to transparency and reproducibility: To foster transparency and encourage further research, we openly share our code and analysis [^1] Source code available at github.com/kahramankostas/GeMID, providing the community with the resources to replicate and build upon our findings.
This paper is organized as follows: Section II reviews related work, Section III covers data selection and feature extraction, Section IV details the process used for selecting and evaluating features, Section V presents model selection and the results of our model generalizability study, Section VI discusses limitations, and Section VII concludes.
II Related Work
The field of IoT DI has been developing in earnest since 2017. Among the pioneering works is IoTSentinel [14], which builds models using 23 features extracted from packet headers. Packet headers, analogous to envelopes for data traveling over a network, contain information for routing and processing the data. By analyzing these headers, various useful features for network tasks can be extracted. Fig. 1 illustrates a sample network packet and lists the header features present in it. Extractable features from packet headers include source/destination IP/MAC addresses, protocol, Time-to-Live (TTL), flag information, port numbers, sequence/acknowledgment numbers, and checksum. A significant contribution of this study was the introduction of  Aalto University IoT device captures [15], one of the most widely used open-access datasets in the IoT DI domain. Subsequent studies, such as IoTSense [9], expanded the feature sets by incorporating payload features (such as payload size and entropy) alongside header-derived features. Concurrently, other research focused on statistics derived from headers, emphasizing the relationship between packets within a specified range (e.g., TTL [8], packet size [16], time [16]).
Figure 1: Raw bytes and headers of a network packet from the Fitbit Aria WiFi enabled weighing device.
In IoTSentinel and IoTSense, features are derived from individual packet headers, but ML models are trained using features collected from multiple packets, known as fingerprints. These fingerprints consist of 12 packets in IoTSentinel and 20 packets in IoTSense. Despite using multiple packets, these fingerprints simply concatenate individual packet features, rather than generating inter-packet features. Alternatively, SysID [10] and IoTDevID [13] create fingerprints based on individual packet features, avoiding the need for merged data features. Obtaining fingerprints from an individual packet addresses the transfer problem that occurs when multiple devices share the same MAC address, a common issue due to the use of MAC addresses in the preassembly of packets [13].
Deriving features from flow data instead of packet headers is also common. A flow refers to a sequence of packets sent from a specific source to a specific destination, considered as part of a single session or connection. The size of the flow can be a predetermined number of packets or packets within a specific time/protocol interval. These features can include basic flow information (source, destination, protocol), time-related metrics (start-end time, duration), packet and byte counts (total, average, max, min sizes), rate metrics (packet, byte, flow rate), flag statistics, inter-arrival times (average, max, min), and flow direction indicators (packets/bytes sent between source and destination). Sivanathan et al. [17] used features obtained from network flows, such as flow volume, flow duration, and average flow rate. They also introduced the widely used UNSW-DI dataset. Many other DI studies have also used flow features [18, 19].
Features can also be obtained from packet headers or flows through sliding windows of various sizes. For example, features such as port ranges, packet size, packet quantity, availability time and inter-arrival time can be subjected to statistical analysis within the window to derive features such as maximum, minimum, mean and standard deviation [20]. This inherently involves features based on multiple packets. However, an issue with features derived from flow/network statistics and windowing methods is that they are likely to capture implicit characteristics of the network environment, rather than just the characteristics of individual devices. In this paper, we argue that this makes them fragile, and demonstrate that they are not a reliable basis for carrying out DI.
The use of raw data is also possible. As shown in Fig. 1, a network packet can be represented as raw bytes. In this approach, the first $n$ bytes of the packet are provided as input to the ML models. Instead of being used explicitly as features, this data is fed into neural network-based learning models, in particular Convolutional Neural Networks (CNNs), where each byte is treated as an image pixel [21, 22, 23, 24, 25]. If the number of bytes exceeds $n$, truncation is applied; if less, padding is used. This method is conceptually similar to the packet header method; but its fields are not strictly separated.
However, using raw data for DI presents other problems that have not yet been adequately addressed. In studies where the entire raw packet data was used [22, 24], this included headers, payloads, and metadata, making it difficult to filter out identifying information such as MAC/IP addresses or string identifiers, both of which can significantly impair model generalizability. Whilst this dependency can be reduced by solely using packet payloads [23, 26], caution is needed because much of today’s internet traffic is encrypted. In practice encrypted payloads are likely to be ineffective for device identification models as the data appears random and opaque, providing no useful information about the contents. Thus, the added complexity of raw data is unlikely to contribute to the model’s performance and may even hinder it.
A more general issue with existing DI studies is the lack of cross-dataset validation, which involves evaluating a model trained on one dataset against a different dataset to assess its generalizability and robustness. While some studies use multiple datasets [11, 12, 13], they typically apply their methods separately to each, training and testing within the same dataset (e.g., training on a portion of Dataset A and testing on the remainder). This approach raises potential concerns about their performance in unseen environments. It also exacerbates the specific issues raised above surrounding the use of statistical features, since models using these features would — in effect — be tested within the same network environment as the one they were trained in. In this study, we demonstrate the importance of cross-dataset validation in developing generalizable models.
III Materials And Methods
III-A Data Selection
DI research typically relies on datasets comprising data collected from real devices. These datasets commonly feature over 20 devices and necessitate prolonged observation periods to gather sufficient data. However, the data collection process for IoT devices entails substantial resource allocation, including personnel and space, making it burdensome. Consequently, many researchers opt to utilize publicly available DI datasets, thereby enhancing reproducibility and transparency and facilitating comparative analysis.
In our study, we use data from two dataset families, UNSW and MonIoTr. Each family consists of multiple datasets that are useful for our model building and evaluation. We use the UNSW datasets for feature and model selection, and the MonIoTr datasets to test the robustness and generalizability of the resulting models.
III-A1 UNSW
For feature and model selection, we utilized the IoT Attack Traces—ACM SOSR 2019 (UNSW-AD) [27] and IoT Traffic Traces—IEEE TMC 2018 (UNSW-DI) [17] datasets. The UNSW-DI dataset is tailored for DI tasks, comprising data from 24 benign devices collected over a 60-day period in 2016. In contrast, the UNSW-AD dataset, originally designed for anomaly detection (AD), contains 27 days of benign data and 17 days of both benign and malicious data from 28 devices collected over a 44-day period in 2018. For the UNSW-AD dataset, we used only the benign data.
Notably, despite both datasets containing most devices in common (see Fig. 2), they were collected at different times, in distinct site environments with varying network characteristics, and for disparate purposes, likely by different users. We leverage these datasets for feature and model selection, employing UNSW-DI as training data and UNSW-AD as test data, and vice versa in different runs/iterations, to enhance model robustness and generalizability.
Each dataset contains a number of sessions with one session per day (UNSW-DI: 60 sessions, UNSW-AD: 27 sessions). The resultant datasets are very large, and some sessions lack certain devices. To address this we selected and merged some sessions to maximize the number of devices available, and then used these merged sessions. Specifically, in the DI dataset, sessions 16-10-03 and 16-11-22 were combined to form S1, while sessions 16-09-29 and 16-11-18 were combined to form S2 (names are in yy-mm-dd date format). In the AD dataset, sessions 18-10-13 and 18-06-14 were combined to form S1, and sessions 18-10-16 and 18-06-11 were combined to form S2.
Figure 2: Devices represented in the UNSW-DI (blue) and UNSW-AD (yellow) datasets, along with their intersection (grey).
Figure 3: Devices represented in the MonIoTr dataset from both UK (blue) and USA (yellow) sites, along with their intersection (grey). Adapted from [28].
III-A2 MonIoTr
After selecting the features and models, to remove any bias, we evaluate their effectiveness on a different dataset, IoT Information Exposure—IMC’19 (MonIoTr) [28]. The MonIoTr dataset consists of 81 devices. An important characteristic of this dataset is that data collection took place at two separate sites, one in the UK (MonIoTr-UK) and the other in the USA (MonIoTr-USA). Of the 81 devices, 26 devices were present in both sites, with the USA site having 47 devices (including 21 unique devices) and the UK site having 34 devices (including 8 unique devices). Fig. 3 shows the distribution of devices by country.
Data collection in these sites spanned 112 hours, covering various scenarios:
{LaTeXdescription}
Data collected during an 8 hour period at night when devices were inactive.
Data collected for 2 minutes immediately after device power-on, without interaction.
Data collected through various interactions, including physical button presses, voice commands, phone app usage on the same network, communication with the device through a phone on a different network utilizing cloud infrastructure, and interaction with the device via Alexa Echo Spot.
Additionally, each device was connected via a VPN to the other site, enabling data collection across both sites for every condition. The same conditions were repeated using VPN tools to establish connections between the sites. For example, idle data was collected by connecting to the UK site from the USA via VPN and vice versa.
Many devices interact heavily with the local network, applications, tools, and cloud services during startup, so we merged power and interaction data into one active class. In a previous study  [29], we observed that when both active and idle data are available, active data is much more effective for DI. Consequently, we excluded idle data from this study.
III-B Feature Extraction
For feature extraction, we used Tshark, a network protocol analyzer, integrated with Python. We focused on features contained in the protocol headers of DNS, HTTP, ICMP, STUN, TCP, UDP, DHCP, EAPOL, IGMP, IP, NTP, and TLS. We eliminated features that consisted of string expressions or contained MAC and IP addresses. After this preliminary elimination, over 300 features remained from 4600 features.
IV Feature Selection
During the feature selection phase, we aimed to identify the most effective features for DI. For this, we used the four UNSW data partitions described in Section III-A1 to guide feature selection. These comprise two sessions (S1 and S2) from each of UNSW-AD and UNSW-DI (referred to hereafter according to their dataset-session number: AD-S1, AD-S2, DI-S1, and DI-S2).
The generalizability of features is assessed in the following three ways, in order of increasing strictness: {LaTeXdescription}
within a partition, which is common in the literature, e.g., within AD-S1.
where generalizability is measured between two partitions of the same dataset, e.g., AD-S1 vs AD-S2.
where generalizability is measured between partitions of different datasets, e.g., AD-S1 vs DI-S2.
IV-A Predictive Features
First, each feature is evaluated individually using each relevant combination of partitions, leading to 16 distinct evaluation contexts (see Fig. 4). Generalizability is assessed by training a decision tree (DT) model using the feature, with DT selected for speed and explainability. The utility of each feature is measured using the kappa metric, which adjusts for chance agreement.
Fig. 5 shows the resulting utility values. Notably, the scores measured under CV are much higher for some features than those measured under DD and SS scores. This indicates that information leaks may be occurring due to train and test folds being taken from the same partition, and suggests that CV is not a reliable basis for selecting features. Hence we only use DD and SS from now on.
A voting system was used for feature selection, with each non-zero kappa value (with a tolerance of ±$\pm 5\%$, so $\geq 0.05$), indicating the feature had a positive contribution, resulting in a vote. Fig. 6 illustrates this system. Features receiving four or more votes from either DD (green) or SS (red) categories, with at least one DD vote, advanced to the next stage. Features not meeting this criterion were eliminated. The requirement for at least one DD vote is based on its greater strictness in assessing generalizability. As a result of voting, we identified 46 of the 332 features as being individually predictive.
Figure 4: Data usage from UNSW datasets across different evaluation contexts. CV, SS, and DD steps are visualized in shades of blue, red, and green, respectively, throughout the paper for clarity. In the nomenclature, CV|AD-S1 denotes cross-validation on the AD dataset session 1. For cases other than CV, the first dataset is used for training and the second for testing. For example, AD-S1|DI-S2 means the first session of the UNSW-AD dataset is used for training and the second session of the UNSW-DI dataset is used for testing.
Figure 5: Comparison of feature utility in UNSW datasets measured using CV (blue) and isolated methods, SS (red) and DD (green). CV tends to overestimate feature utility, with higher scores for many attributes. SS and DD produce more realistic evaluations. The discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models.
Figure 6: Features identified as being predictive within one or more evaluation contexts (CV: blue, SS: red, DD: green), based on non-zero kappa scores of DT models. Note that the CV votes are not used to select features, but are shown here for completeness.
Figure 7: List of intersecting features identified across eight dataset versus dataset (DD) cases using Genetic Algorithm (GA). Feature names highlighted in dark green were selected for model building after further post hoc analysis.
IV-B Feature Interactions
We then considered feature interactions, using a wrapper method, specifically a Genetic Algorithm (GA), to find the optimal combination of individually predictive features which we validated across multiple datasets. The global search characteristic of a GA tends to find good solutions, and avoiding exhaustive search enables us to obtain potential feature sets within an acceptable timeframe. However, in general there is no guarantee of a GA finding an optimal subset, meaning it may include uninformative features in the feature set. Additionally, GA-selected features may develop a bias towards the data used for selection, potentially reducing their performance on other datasets.
To address this, all 46 features that passed individual voting were combined into a single feature set for selection using a genetic algorithm (GA). A decision tree (DT) model was employed for evaluation, with the fitness of each generation assessed using the F1 score across 8 additional DD datasets (see Fig. 4). This external validation provides feedback to the GA, ensuring that feature selection is based on success across different datasets and not dependent on only one. For each of the eight DD cases, the GA produced slightly different feature sets, so we then identified the intersections between them. Fig. 7 depicts this, showing that some features (e.g., tcp.window_size, tcp.ack) were chosen multiple times, some (e.g., tcp.seq, tcp.flags.ack) only once.
IV-C Deriving a Final Feature Set
To identify the most effective combination of features from the GA results, we conducted two post hoc analyses. First, the feature set derived from each DD case (each GA run) is reevaluated on each other DD case. F1 scores of the resulting DT models are presented in the left part of Fig. 8. These show how well the feature sets from each GA run generalize.
The second analysis uses a voting mechanism based on intersections of the eight GA feature sets (Fig. 7), with a feature receiving votes proportional to how often it appears. Features are then grouped based on thresholds, i.e., whether they appear in 2 or more feature sets (Vote+2), 3 or more (Vote+3) etc., and DT models are trained using these feature groups. The results are shown on the right side of Fig 8.
From the mean F1 scores shown in the final row of Fig. 8, it can be seen that feature sets based on grouping generally lead to better performance, at least for voting thresholds up to 4. For larger thresholds, the feature sets become very small, explaining the lower scores. Generally, the feature sets from single runs lead to models with lower F1 scores, suggesting that voting adds more robustness. Consequently, we use the Vote+3 feature set — which has the highest F1 score — in the next section. The selected features are highlighted in Fig. 7.
Figure 8: Comparison of feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all DD data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
V Model Evaluation
Next, we construct and evaluate ML models using the selected features and compare their performance against established baselines using an independent dataset.
V-A Model Selection
We experimented with various ML modeling approaches commonly used in DI to identify the most effective one for this purpose, in terms of both predictive success and inference time — the latter being an important consideration when scanning network packets. The models we considered were Logistic Regression (LR), Decision Trees (DT), Naive Bayes (NB), Support Vector Machines (SVM), Random Forest (RF), Extreme Gradient Boosting (XGB) Multi-Layer Perceptron (MLP), K-Nearest Neighbors (KNN), Convolutional Neural Networks (CNN), Long Short-Term Memory (LSTM), and Bidirectional Encoder Representations from Transformers (BERT).
We used random search for hyperparameter optimization and applied each model to the DD datasets — see Supplementary Material (SM) Section 4.4 for details. The F1 scores and average inference times are shown in Fig. 9. From this, it is clear that the most successful models are RF and XGB, with F1 scores of 0.799 and 0.780, respectively. While their predictive performance is similar, RF runs about 6 times faster than XGB, so we chose this as our preferred model. Notably, DT has the fastest inference time, but its predictive performance lags behind RF by about 10 percentage points.
Figure 9: F1 scores and average inference times of various ML algorithms applied to DD datasets. RF and XGB demonstrate the highest F1 scores of 0.799 and 0.780, respectively, while DT show the fastest inference time.
V-B Evaluating Model Generalizability
In this section, we measure the generalizability of our method (GeMID) and other methods using the MonIoTr dataset, using the US, US-VPN, UK, and UK-VPN partitions of the MonIoTr dataset within three different evaluation contexts. In CV, we carry out cross-validation within a single partition. In SS, we train models using non-VPN data and test them using VPN data, and vice versa. In DD, we use data collected from one geographical site to train models, and data collected from the other geographical site to test them.
As comparison baselines, we use flow-based features generated by CICFlowmeter [30] and window-based features generated by Kitsune [31]. These are well-known tools that were used in previous DI studies. To ensure a fair comparison, feature and model selection were repeated independently (see SM Section I and II for results). We also compare against IoTDevID [13], another packet header-based DI method that aimed to build generalizable DI models and demonstrated improved performance over alternative methods.
TABLE I: Comparison of DI methods on MonIoTr dataset. F1 scores.
Dataset GeMID CICFlwM IoTDevID Kitsune
CV UK 0.960 0.857 0.900 0.931
UKVPN 0.965 0.853 0.908 0.924
US 0.965 0.904 0.910 0.903
USVPN 0.954 0.904 0.923 0.944
Mean 0.961 0.879 0.910 0.925
SS UK|UKVPN 0.895 0.644 0.824 0.741
UKVPN|UK 0.877 0.555 0.802 0.710
US|USVPN 0.861 0.606 0.821 0.712
USVPN|US 0.894 0.616 0.839 0.687
Mean 0.882 0.605 0.822 0.712
DD UK|US 0.760 0.475 0.703 0.464
UK|USVPN 0.769 0.490 0.666 0.435
UKVPN|US 0.748 0.516 0.654 0.392
UKVPN|USVPN 0.746 0.454 0.709 0.444
US|UK 0.778 0.454 0.752 0.541
US|UKVPN 0.780 0.517 0.694 0.455
USVPN|UK 0.811 0.515 0.710 0.439
USVPN|UKVPN 0.816 0.464 0.731 0.521
Mean 0.776 0.486 0.702 0.461
The results are shown in Table I. All methods achieved high scores for CV, close to or above 0.90. However, a decrease is observed in all methods for SS and DD, with this decrease being particularly dramatic for the statistics-based methods (CICFlowMeter and Kitsune). For DD, GeMID maintains 81% of its CV F1 score and IoTDevID 77%, compared to only 55% for CIC and 50% for Kitsune. This significant drop supports our claim that flow or window-based statistical methods, which focus on relationships between devices within the network rather than individual device activities, are not suitable for building generalizable DI models due to their poor transferability across heterogeneous network environments.
The two methods — GeMID and IoTDevID — based on packet headers both perform significantly better, with GeMID outperforming IoTDevID within all evaluation contexts. Table II shows the per-device results for GeMID within the DD evaluation context. It can be seen that most of the devices are identified with a high level of accuracy. A notable exception is the echo family of devices, which, presumably due to underlying similarities, are easily misclassified with each other (see SM Figs. 12 and 13 for confusion matrices). There is also a degree of misclassification between devices from different manufacturers that perform the same function, e.g., between roku-tv and samsungtv-wired, and between yi-camera and wansview-cam-wired. A third category of poor performance is due to data paucity, e.g., for sousvide, there are only 70 US samples, compared to 15,000 for UK. The challenges in classifying these devices are not specific to GeMID but are also observed in other methods (IoTDevID, CICFlowMeter, and Kitsune). These methods also encounter similar difficulties due to the inherent complexity of distinguishing between devices with similar characteristics.
TABLE II: F1 scores per device in all 8 DD cases for MonIoTr data, with green for higher and red for lower success.
Devices Train→ UK UK UKVPN UKVPN US US USVPN USVPN
Test→ US USVPN US USVPN UK UKVPN UK UKVPN
appletv 0.868 0.879 0.851 0.830 0.658 0.858 0.614 0.802
blink-camera 0.963 0.983 0.952 0.985 0.951 0.929 0.945 0.932
blink-security-hub 0.909 0.889 0.498 0.775 0.259 0.213 0.649 0.599
echodot 0.161 0.093 0.106 0.053 0.521 0.083 0.575 0.313
echoplus 0.263 0.387 0.430 0.242 0.402 0.590 0.479 0.413
echospot 0.629 0.542 0.633 0.532 0.704 0.798 0.760 0.739
firetv 0.629 0.695 0.646 0.694 0.638 0.809 0.651 0.805
google-home-mini 0.806 0.885 0.761 0.873 0.873 0.895 0.909 0.903
insteon-hub 0.909 0.842 0.648 0.537 0.806 0.712 0.958 0.962
lightify-hub 0.951 0.941 0.946 0.936 0.839 0.924 0.969 0.979
magichome-strip 0.837 0.857 0.905 0.905 0.977 0.945 0.978 0.952
nest-tstat 0.879 0.803 0.870 0.866 0.959 0.945 0.957 0.956
ring-doorbell 0.997 0.998 0.997 0.999 0.996 0.991 0.995 0.991
roku-tv 0.337 0.490 0.482 0.507 0.692 0.693 0.448 0.563
samsungtv-wired 0.834 0.848 0.941 0.908 0.808 0.820 0.659 0.645
sengled-hub 0.670 0.758 0.954 0.922 0.525 0.455 0.533 0.552
smartthings-hub 0.853 0.880 0.910 0.907 0.932 0.904 0.934 0.911
sousvide 0.422 0.394 0.128 0.092 0.992 0.982 0.997 0.994
t-philips-hub 0.987 0.988 0.987 0.987 0.956 0.826 0.977 0.971
t-wemo-plug 0.935 0.844 0.938 0.841 0.957 0.979 0.957 0.982
tplink-bulb 0.941 0.977 0.852 0.946 0.904 0.857 0.994 0.987
tplink-plug 0.783 0.905 0.836 0.929 0.832 0.855 0.808 0.886
wansview-cam-wired 0.875 0.822 0.876 0.841 0.914 0.917 0.908 0.913
xiaomi-hub 0.825 0.879 0.830 0.870 0.753 0.924 0.986 0.989
yi-camera 0.729 0.654 0.719 0.673 0.596 0.589 0.646 0.657
Mean accuracy 0.857 0.825 0.861 0.830 0.893 0.888 0.898 0.889
Mean F1 score 0.760 0.769 0.748 0.746 0.778 0.780 0.811 0.816
V-C Feature Usage
GeMID and IoTDevID demonstrate better generalization in model performance compared to those that use statistical features. Of these two, GeMID is approximately 8 percentage points higher in accuracy. Given that the model building stages are similar, this is presumably due in part to a more rigorous feature selection process, using two data sets rather than just one to ensure that the selected features are generalizable.
Notably, IoTDevID’s feature set is rich in features related to BOOTP, DNS, and EAPOL protocols, whereas GeMID lacks these. This may be due to a dataset-specific bias, owing to the reliance on a single dataset — Aalto — for IoTDevID’s feature selection process. This is supported by the observation that while GeMID outperforms IoTDevID in experiments using the UNSW and MonIoTr datasets, IoTDevID shows markedly better performance when applied to the Aalto dataset [13] (see SM Table VIII).
However, it is worth noting that both share the features dstport_class, ip.flags.df, ip.len, ip.ttl, and tcp.window_size, suggesting that these are particularly important for DI.
Another reason for the difference in performance may be the larger initial pool of extracted features in GeMID (332 features) compared to IoTDevID (112 features). Although both studies generally use the same protocols, GeMID captures more sub-features within each protocol, providing finer-grained information in the feature set. In terms of selected features, GeMID offers more comprehensive protocol coverage, starting with a richer set of HTTP features and extending to broader coverage of TCP and UDP protocols. This includes features such as sequence numbers, timing metrics, and window size information, which enhance its ability to capture more detailed network behavior (see Fig. 7). Unlike in the IoTDevID study, we retained features containing session-based identifiers. Whilst these can lead to over-optimistic metric scores [32], this is only the case when sessions are split between test and train sets, e.g., when using CV. When this is controlled for using stricter evaluation contexts (as in this study), the inclusion of these features can enhance model performance by capturing meaningful relationships between packets.
In practice, the accuracy of packet-level classification methods can be improved by aggregating predictions from individual packets. The IoTDevID study introduced an aggregation algorithm that enhances performance by combining multiple packet-level predictions. This approach is particularly effective for packet-level methods such as GeMID, IoTDevID, and Kitsune (despite Kitsune focusing on the relationship between the window system and packets, as it still relies on packet-based labels). In contrast, CICFlowMeter operates at the flow level, where such aggregation is not applicable. Our results in SM Tables VI and VII show the impact of this approach, with GeMID achieving a mean F1 score of 0.892 in the DD evaluation context.
VI Limitations
The primary limitation of this work is the availability and quality of public datasets. Although we used well-regarded datasets and cross-dataset validation, there will always be a limit in terms of the diversity, representativeness and sampling biases of any dataset.
Future work in DI would benefit from larger datasets designed specifically for this purpose, containing device samples that reflect realistic usage situations. Whilst we already consider devices situated in two different network environments, collecting data from the same devices operating in a broader range of networks would inevitably provide a broader perspective on generalizability. There is also a need for more data from non-IP protocols such as ZigBee and Z-Wave to test how well DI models of non-IP devices generalize.
Whilst this study (in common with previous studies) focuses on benign data, future DI studies would also benefit from using attack or malicious data, since this would be more representative of the actual environments within which DI models are deployed.
VII Conclusions
Our work presents a novel approach to improving the generalizability of IoT DI models. Using a two-stage framework involving feature and algorithm selection followed by cross-dataset validation, we show that models trained on datasets in one environment can effectively identify devices in another environment. This addresses an important gap in existing methodologies, which often do not validate across different datasets.
Comparative analysis of DI methods showed that packet header-based features performed better than network statistics and window methods. Packet features offer a more consistent and reliable basis for model training because they are less influenced by network-specific variables. This finding highlights the importance of selecting features that inherently capture device-specific attributes.
Our results also highlight the limitations of relying only on single datasets for model training and validation. Such approaches run the risk of overfitting and may fail to generalize to different network environments. In contrast, our methodology provides robustness and adaptability, which are crucial for practical IoT security applications.